Bowie Sporting Goods manufactures sleeping bags. The manufacturing standards per sleeping bag, based on 5,000 sleeping bags per month, are as follows: Direct material of 4.00 yards at $5.75 per yard Direct labor of 2.50 hours at $16.00 per hour Overhead applied per sleeping bag at $16.00 In the month of April, the company actually produced 5,200 sleeping bags using 27,300 yards of material at a cost of $5.50 per yard. The labor used was 11,700 hours at an average rate of $19.50 per hour. The actual overhead spending was $96,200. Determine the labor rate variance and round to the nearest whole dollar. Enter a favorable variance as a negative number. Enter an unfavorable variance as a positive number.
Answer #1

Labor rate variance. = (SR-AR) * AH

= (16 - 19.50) * 11,700 Hours

= - 3.50 * 11,700

= 40,950 Unfavorable

Comment if you face any issues